# Graph Estimation Based on Neighborhood Selection for Matrix-variate Data

### v1

@@ 1 Introduction
We conclude this section by briefly introducing the notations used throughout this paper. Let $[p]=\{1,2,\ldots,p\}$ for a positive integer $p$. Denote by $|S|$ the cardinality of a set $S$. For $k\in[p]$, define $(k)=\{1,\ldots,k-1,k+1,\ldots,p\}$. For a vector $x\in\mathbb{R}^{p}$ and $k\in[p]$, denote a leave-one-out vector by $x_{(k)}=(x_{1},\ldots,x_{k-1},x_{k+1}\ldots,x_{p})^{\top}\in\mathbb{R}^{p-1}$. For $p\times q$ matrix $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, $\boldsymbol{X}_{(k)}$ is a $(p-1)\times q$ matrix where $X_{k}$ is eliminated. Moreover, for $S_{1}\subset[p]$ and $S_{2}\subset[q]$, let $\boldsymbol{X}_{S_{1}S_{2}}$ be the sub-matrix consisting of rows and columns indexed by $S_{1}$ and $S_{2}$, respectively. For a singleton set $S=\{s\}$, we use $s$ instead of $S$ to denote the sub-matrix. For example, $X_{s}$ denotes the $s$-th row of $\boldsymbol{X}$.
@@ 3 Theoretical results
[…]
@@ 5 Real data analysis
Figure 5: The estimated column-wise graph of temporal measurements using matrixNS with $\lambda=4.0$. Nodes (125612561–2561 – 256) correspond to individual measurement epochs.
@@ A.1 Proof of Assertion (i) of Theorem 1
To prove (i), we employ the primal-dual witness argument of Wainwright, (2019). Specifically, we seek an optimal primal-dual pair $(\hat{\theta},\hat{z})$ that satisfies $\mathbb{X}_{(a)}^{\top}(\mathbb{X}_{(a)}\hat{\theta}-\mathbb{X}_{a})/n+\lambda\hat{z}=0$. A straightforward calculation shows
@@
[…]
@@
Since $\tilde{\boldsymbol{W}}_{S^{c}}^{\top}$ and $\epsilon$ are independent, $T_{S^{c}}|\epsilon$ follows the centered normal distribution. Denote its $k$th component by $t_{k}:=e_{k}^{\top}T_{S^{c}}=\tilde{W}_{k}^{\top}(\text{I}-\boldsymbol{\Pi}_{S})\epsilon/(nq\lambda)$ On $E_{2a}$, its conditional variance satisfies
@@ A.1.1 Proof of Fact 1
We now show $E_{1a}^{c}$ holds with high probability. By Lemma 1 with $t_{0}=\lambda_{\min}^{U}v_{\text{avg}}/(2C||\boldsymbol{U}||_{2})$, one obtains
@@
Under the mild assumption $n(2d_{\max}/q+1)>||\boldsymbol{V}||_{2}^{2}$, then all the three terms
$\frac{a_{1}^{3/2}/a_{3}^{1/2}}{a_{2}^{3}/a_{3}^{2}}=\left(\dfrac{n(2d_{\max}/q+1)}{||\boldsymbol{V}||_{2}^{2}}\right)^{3/4},\ \frac{a_{1}^{2}/a_{2}}{a_{1}^{3/2}/a_{3}^{1/2}}=\left(\dfrac{n(2d_{\max}/q+1)}{||\boldsymbol{V}||_{2}^{2}}\right)^{1/4},\ \frac{a_{1}^{2}/a_{2}}{a_{2}^{3}/a_{3}^{2}}=\dfrac{n(2d_{\max}/q+1)}{||\boldsymbol{V}||_{2}^{2}}$
@@
$n\!\left(\frac{2d_{\max}}{q}\!+1\right)\!>\!||\boldsymbol{V}||_{2}^{2},\ \ \frac{n}{\log p}\geq\beta\!\left(\frac{2d_{\max}}{q}+1\right)\!\frac{||\boldsymbol{V}||_{2}^{2}}{t_{0}^{2}},\ \ \frac{q^{2}}{n}\geq\left(\frac{2d_{\max}}{q}\!+1\right)^{-3}\!\frac{t_{0}^{4}}{||\boldsymbol{V}||_{2}^{2}}.$ (15)
@@
$n\left(\frac{2d_{\max}}{q}\!+1\right)\!>\!||\boldsymbol{V}||_{2}^{2},\ \ \frac{nq}{(\log p)^{3/2}}\geq\beta^{3/2}\frac{||\boldsymbol{V}||_{2}^{2}}{t_{0}},\ \ \frac{n}{q^{2}}\geq\left(\frac{2d_{\max}}{q}\!+1\right)^{3}\!\frac{||\boldsymbol{V}||_{2}^{2}}{t_{0}^{4}}.$ (16)
@@ B.1 Lemma 1
$\mathbb{P}\left(\bigg{|}\bigg{|}\sum_{i=1}^{n}\big{\{}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}\text{I}_{A_{i}}-\mathbb{E}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}\text{I}_{A_{i}}\big{\}}\bigg{|}\bigg{|}_{2}\geq n\delta\right)\leq 2m\exp\left\{-\dfrac{1}{4}\min\bigg{\{}\dfrac{n\delta^{2}}{(m+1)\|\boldsymbol{V}\|_{F}^{2}+\text{tr}(\boldsymbol{V})^{2}},\dfrac{3n\delta}{2\|\boldsymbol{V}\|_{2}^{2}K}\bigg{\}}\right\}.$
@@
$\frac{1}{n}\bigg{|}\bigg{|}\sum_{i=1}^{n}\big{\{}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}\text{I}_{A_{i}}-\mathbb{E}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}\text{I}_{A_{i}}\big{\}}\bigg{|}\bigg{|}_{2}\geq 4\max\bigg{\{}\sqrt{\dfrac{\Big{(}(m+1)\|\boldsymbol{V}\|_{F}^{2}+\text{tr}(\boldsymbol{V})^{2}\Big{)}u}{n}},\dfrac{2\|\boldsymbol{V}\|_{2}^{2}Ku}{3n}\bigg{\}}$
@@
$\bigg{|}\bigg{|}\mathbb{E}\boldsymbol{Z}^{(1)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(1)}_{S})^{\top}\text{I}_{A_{1}^{c}}\bigg{|}\bigg{|}_{2}$ $=\max\limits_{x:\|x\|_{2}=1}\mathbb{E}(x^{\top}\boldsymbol{Z}^{(1)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(1)}_{S})^{\top}x)\cdot\text{I}_{A_{1}^{c}}$
@@
Here $S$ is treated as fixed, so the $q$–dimension plays the role of sample size. Setting $\delta=\sqrt{2(u+\log n)/q}$ for $u>0$ gives
@@
$\bigg{|}\bigg{|}\mathbb{E}\boldsymbol{Z}^{(1)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(1)}_{S})^{\top}\text{I}_{A_{1}^{c}}\bigg{|}\bigg{|}_{2}\leq\sqrt{\text{tr}(\boldsymbol{V})^{2}+2\|\boldsymbol{V}\|_{F}^{2}}\ e^{-u/2}\leq\sqrt{\dfrac{\text{tr}(\boldsymbol{V})^{2}+2\|\boldsymbol{V}\|_{F}^{2}}{n}}.$
@@
$\mathbb{P}\left[\bigg{|}\bigg{|}\frac{1}{n}\sum_{i=1}^{n}\big{\{}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}-\mathbb{E}\boldsymbol{Z}^{(i)}_{S}\boldsymbol{V}(\boldsymbol{Z}^{(i)}_{S})^{\top}\big{\}}\bigg{|}\bigg{|}_{2}\geq t_{1}+t_{2}\right]$

### v2

@@ 1 Introduction
We conclude this section by briefly introducing the notations used throughout this paper. Let $p]=\{1,2,\ldots,p\}$ for a positive integer $p$. Denote by $|S|$ the cardinality of a set $S$. For $k\in[p]$, define $(k)=\{1,\ldots,k-1,k+1,\ldots,p\}$. For a vector $x\in\mathbb{R}^{p}$ and $k\in[p]$, denote a leave-one-out vector by $x_{(k)}=(x_{1},\ldots,x_{k-1},x_{k+1}\ldots,x_{p})^{\top}\in\mathbb{R}^{p-1}$. For $p\times q$ matrix $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, $\boldsymbol{X}_{(k)}$ is a $(p-1)\times q$ matrix where $X_{k}$ is eliminated. Moreover, for $S_{1}\subset[p]$ and $S_{2}\subset[q]$, let $\boldsymbol{X}_{S_{1}S_{2}}$ be the sub-matrix consisting of rows and columns indexed by $S_{1}$ and $S_{2}$, respectively. For a singleton set $S=\{s\}$, we use $s$ instead of $S$ to denote the sub-matrix. For example, $X_{s}$ denotes the $s$-th row of $\boldsymbol{X}$.
@@ 3 Theoretical results
Based on the matrix neighborhood selection (4), estimate the edge set $\hat{E}$ by either $\hat{E}_{\text{AND}}$ or $\hat{E}_{\text{OR}}$ as in Section 2.2. Then, with probability at least $1-\max(3d_{\max}+6,\ 5d_{\max})/p^{\beta-1}$, the estimated edge set $\hat{E}$ successfully recovers the true set $E_{U}=\{(a,b)\in[p]\times[p]:(\boldsymbol{U}^{-1})_{ab}\neq 0\}$ according to the following criteria:
@@ 5 Real data analysis
Figure 5: The estimated column-wise graph of temporal measurements using matrixNS with $\lambda=4.0$. Nodes (12561–256) correspond to individual measurement epochs.
@@ A.1 Proof of Assertion (i) of Theorem 1
[…]
@@
Thus we can write $\epsilon^{(i)}\overset{d}{=}{\boldsymbol{V}^{*}}^{1/2}Z^{(i)}$ with $Z^{(i)}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}_{q}(0,I_{q})$. Let $\tilde{Z}=(Z^{(1)},\dots,Z^{(n)})^{\top}$ and $\tilde{\boldsymbol{V}}^{*}=\mathrm{diag}_{n}(\boldsymbol{V}^{*})$, then the squared norm becomes
@@
Since $\tilde{\boldsymbol{W}}_{S^{c}}^{\top}$ and $\epsilon$ are independent, $T_{S^{c}}|\epsilon$ follows the centered normal distribution. Denote its $k$th component by $t_{k}:=e_{k}^{\top}T_{S^{c}}=\tilde{W}_{k}^{\top}(\text{I}-\boldsymbol{\Pi}_{S})\epsilon/(nq\lambda)$ . On $E_{2a}$, its conditional variance satisfies
@@ A.1.1 Proof of Fact 1
[…]
@@
[…]
@@
[…]
@@
[…]
@@ B.1 Lemma 1
[…]
$\qquad\qquad\qquad\qquad\qquad\leq 2m\exp\left\{-\dfrac{1}{4}\min\bigg{\{}\dfrac{n\delta^{2}}{(m+1)\|\boldsymbol{V}\|_{F}^{2}+\text{tr}(\boldsymbol{V})^{2}},\dfrac{3n\delta}{2\|\boldsymbol{V}\|_{2}^{2}K}\bigg{\}}\right\}.$
@@
[…]
$\qquad\qquad\qquad\qquad\qquad\geq 4\max\bigg{\{}\sqrt{\dfrac{\Big{(}(m+1)\|\boldsymbol{V}\|_{F}^{2}+\text{tr}(\boldsymbol{V})^{2}\Big{)}u}{n}},\dfrac{2\|\boldsymbol{V}\|_{2}^{2}Ku}{3n}\bigg{\}}$
@@
[…]
@@
Here $S$ is treated as fixed, so the qq–dimension plays the role of sample size. Setting $\delta=\sqrt{2(u+\log n)/q}$ for $u>0$ gives
@@
[…]